Papa Roach Company uses the percentage of receivables method for recording bad debts expense. The accounts receivable balance is $200,000 and credit sales are $1,000,000. Management estimates that 5% of accounts receivable will be uncollectible. Prepare the year-end journal entry that Papa Roach Company will make if the Allowance for Doubtful Accounts has an existing credit balance of $2,000

Prepare the year-end journal entry.
Answer #1
Explaination :
Estimated uncollectible account at end : 2,00,000*5% = $10,000
Bad debt expense to be recorded =estimated uncollectible - balance in allowance account (cr)
           = 10,000 - 2000 = $8,000
Journal entry :
Accounts title and explaination Debit Credit
Bad debt expense A/c                           Dr. $8,000.00
    To Allowance for doubtful accounts $8,000.00